You are the principal financial analysts of the Gadget Division of The FGM Corporation, the

largest multinational automobile manufacturer in the world. You are asked to evaluate a project

proposal regarding the production of a new voice-activated electronic device – Universal

Direction Assistance (UDA). This upgradeable device, which incorporates the most advanced

computer and satellite wireless technology, provides directions and real time traffic reports that

guide automobile drivers in choosing the preferred route to their destinations. This device can be

used in any country with specialized software that contains local geographical information and

real time traffic report (where technology is available) that are translated into the chosen

language of the driver. UDA will be sold as an option for the FGM cars and trucks. A

comprehensive market analysis on the potential demand for this device was conducted last year

at a cost of $10M.

According to the results of the market analysis, the expected annual sale volumes of UDA are

2.3M units for the first two years, and drop to 2.1M units for the final two years of this 4-year

project. Unit prices are expected to be $650 for the first two years, and then fall to $550, due to

the introduction of similar devices by competitors, afterwards. Unit production costs are

estimated at $520 in the first year of the project. The growth rate for unit production costs are

expected to be 4% per year over the remaining life of the project.

In addition, the implementation of the project demands current assets to be set at 12% of the

annual sale revenues, and current liabilities to be set at 8% of the annual production costs.

Besides, the introduction of UDA will increase the sales volume of cars and trucks that leads to

an increase in the annual after-tax cash flow of FGM by $21M.

The production line for UDA will be set up in a vacant plant that was built by FGM at a cost of

$30M twenty years ago. This fully depreciated plant has a current market value of $20M, and is

expected to be sold at the termination of this project for $12M in four years. The machinery for

producing UDA has an invoice price of $120M, and its modification costs another $15M. The

machinery has an economic life of 4 years, and is classified in the MACR 3-year class for

depreciation purpose. The machinery is expected to have zero salvage value at the termination

of the project.

The cost of capital (or discount rate) for this project is assumed to be 15%. The estimated

marginal tax rate of The FGM Corporation is 21%.

2

After reading the case, please answer the following questions:

1. Set up Excel model to estimate the incremental cash flows to be generated by the project over

the life of the project.

2. Based on the cost of capital of the project, please calculate NPV and IRR of the project.
Answer #1
Computation of Earnings after tax
Year 1 Year 2 Year 3 Year 4 Total
Sales Value 1,495,000,000.00 1,495,000,000.00 1,155,000,000.00 1,155,000,000.00 5,300,000,000.00
Production Costs 1,196,000,000.00 1,243,840,000.00 1,181,103,000.00 1,228,332,000.00 4,849,275,000.00
Contribution     299,000,000.00     251,160,000.00      (26,103,000.00)      (73,332,000.00)     450,725,000.00
Depreciation       44,995,500.00       60,007,500.00       19,993,500.00       10,003,500.00     135,000,000.00
Net Operating Income     254,004,500.00     191,152,500.00      (46,096,500.00)      (83,335,500.00)     315,725,000.00
Tax @ 21%       53,340,945.00       40,142,025.00                        -                          -         93,482,970.00
Earnings after tax     200,663,555.00     151,010,475.00      (46,096,500.00)      (83,335,500.00)     222,242,030.00
Cash Flow after tax Year 1 Year 2 Year 3 Year 4 Total
Earnings after tax     200,663,555.00     151,010,475.00      (46,096,500.00)      (83,335,500.00)     222,242,030.00
Add: Depreciation       44,995,500.00       60,007,500.00       19,993,500.00       10,003,500.00     135,000,000.00
Cash Flow after tax     245,659,055.00     211,017,975.00      (26,103,000.00)      (73,332,000.00)     357,242,030.00
Working Capital Year 1 Year 2 Year 3 Year 4 Total
Current Assets     179,400,000.00     179,400,000.00     138,600,000.00     138,600,000.00     636,000,000.00
Current Liabilities       95,680,000.00       99,507,200.00       94,488,240.00       98,266,560.00     387,942,000.00
Working Capital Investments       83,720,000.00       79,892,800.00       44,111,760.00       40,333,440.00     248,058,000.00
Computation of Incremental Cash Flow
Year 1 Year 2 Year 3 Year 4 Total
Cash inflow from increased sales volume of cars & trucks       21,000,000.00       21,000,000.00       21,000,000.00       21,000,000.00       84,000,000.00
Cash inflow after tax     245,659,055.00     211,017,975.00      (26,103,000.00)      (73,332,000.00)     357,242,030.00
Cash Outflow against
investment in Working Capital
     (83,720,000.00)      (79,892,800.00)      (44,111,760.00)      (40,333,440.00)    (248,058,000.00)
Cash inflow from sale of Plant       12,000,000.00       12,000,000.00
Total Cash inflow     182,939,055.00     152,125,175.00      (49,214,760.00)      (80,665,440.00)     205,184,030.00
Less:
Cash Outflow against purchase of Machinery                        -                          -                          -                          -      (120,000,000.00)
Cash Outflow against modification of Machinery                        -                          -                          -                          -        (15,000,000.00)
Value of Plant      (20,000,000.00)
Net incremental cash flow from the Project     255,368,060.00
Computation of NPV of the Project
Year Net Cash Flow PV Factor @ 15% PV of Cash flow
Year 1     182,939,055.00                     0.87     159,156,978.00
Year 2     152,125,175.00                     0.76     115,615,133.00
Year 3      (49,214,760.00)                     0.66      (32,481,742.00)
Year 4      (80,665,440.00)                     0.57      (45,979,301.00)
Net Present Value of Cash Flow     196,311,068.00
Less: Initial investment    (155,000,000.00)
Net Present Value of Project       41,311,068.00
Computation of IRR of the project
At Internal rate of return,the NPV of the project is zero
Hence, at IRR
(Net Cash flow for the period/(1+IRR) ^ t) - Total Initial Investments = 0
where, IRR is the internal rate of return
and t = the number of years
Net Cash flow for the period     205,184,030.00
Initial Investment     155,000,000.00
putting the values in the formula
=( 205,184,030/((1+IRR) ^ 4)) - 155,000,000 = 0
155,000,000 = 205,184,030/((1+IRR) ^ 4)
= 155,000,000 * ((1+IRR) ^ 4) = 205,184,030
= ((1+IRR) ^ 4) = 205,184,030/155,000,000
= ((1+IRR) ^ 4) = 1.3238
(1+IRR) = (1.3238) ^ (1/4)
1+IRR = 1.0726
IRR = 1.0726 - 1
IRR = 0.726
IRR = 7.26%
Working Notes
Computation of Sales Value
Year 1 Year 2 Year 3 Year 4 Total
Sales Volume         2,300,000.00         2,300,000.00         2,100,000.00         2,100,000.00         8,800,000.00
Unit Price                  650.00                  650.00                  550.00                  550.00
Sales Value 1,495,000,000.00 1,495,000,000.00 1,155,000,000.00 1,155,000,000.00 5,300,000,000.00
Computation of Production Costs
Year 1 Year 2 Year 3 Year 4 Total
Sales Volume         2,300,000.00         2,300,000.00         2,100,000.00         2,100,000.00         8,800,000.00
Unit Cost                  520.00                  540.80                  562.43                  584.92
Production Costs 1,196,000,000.00 1,243,840,000.00 1,181,103,000.00 1,228,332,000.00 4,849,275,000.00
Production Costs increases by 4% year on year
Year Unit Cost Increase
Year 1                  520.00                   20.80
Year 2                  540.80                   21.63
Year 3                  562.43                   22.50
Year 4                  584.93                   23.40
Computation of Current Assets & Liabilities
Year 1 Year 2 Year 3 Year 4 Total
Sales 1,495,000,000.00 1,495,000,000.00 1,155,000,000.00 1,155,000,000.00 5,300,000,000.00
Current assets @ 12% of
Sales
    179,400,000.00     179,400,000.00     138,600,000.00     138,600,000.00     636,000,000.00
Production Costs 1,196,000,000.00 1,243,840,000.00 1,181,103,000.00 1,228,332,000.00 4,849,275,000.00
Current liabilities @ 8% of
Production Costs
      95,680,000.00       99,507,200.00       94,488,240.00       98,266,560.00     387,942,000.00
Computation of Depreciation for MACRS 3 year asset class
Invoice Value of Machinery     120,000,000.00
Cost of Modification       15,000,000.00
Total Cost of Machinery     135,000,000.00
Less: Salvage Value                        -  
Depreciable Value     135,000,000.00
MACRS depreciation schedule for 3 year asset class
Year Asset Value MACRS Depreciation
Rate
Depreciation
    135,000,000.00
Year 1 33.33%       44,995,500.00
Year 2 44.45%       60,007,500.00
Year 3 14.81%       19,993,500.00
Year 4 7.41%       10,003,500.00